OpenStudy (anonymous):

Someone rolls a fair six-sided die and you win points equal to the number shown. What is the expected number of points after one roll?

OpenStudy (anonymous):

3.5 would be the answer

OpenStudy (anonymous):

do 1 * 1/6 + 2 * 1/6 + 3 * 1/6 + 4 * 1/6 + 5 * 1/6 + 6 * 1/6 and that equals 3.5 because its the probability for each roll when can be anything but the official number is 3.5 Do you understand?

OpenStudy (anonymous):

thanks ilikephysics2

OpenStudy (anonymous):

Do you understand though what i did

OpenStudy (unklerhaukus):

\[\langle x\rangle=\sum\limits_{n=1}^{6}xP(x)=\frac16\sum\limits_{n=1}^{6}x\]\[=\frac16\left(1+2+3+4+5+6\right)\]\[=\frac{21}6\]\[=\frac{3\times6+3}{6}\]\[=3+\frac 36\]\[=\]

OpenStudy (unklerhaukus):

i dont understand your first post @ilikephysics2

OpenStudy (anonymous):

just ignore it haha my keepboard messed up